y=x-(x"2)/2+(x"3)/3-(x"4)/4+......PROOF THAT,,[ (1+x)dy/dx=1]

Grade:12

1 Answers

rishabh
7 years ago
there are two ways..... 1> the given function is ln(1+x)=y  (by taylor series) …
2> differentiate on both sides then use.... 1 – x + x^2 …..=1/(1+x) by infinte gp formula...hope it helps